OpenStudy (anonymous):

If an investor purchased 50 shares of IBM stock at $70 per share, 110 shares at $110 per share, 130 shares at $135 per share, and another 75 shares at $145 per share, what is the mean cost per share? Round your answer to 2 decimal places.

OpenStudy (anonymous):

It will be total cost divided by total number of shares. The cost is: \[50\times $70+110 \times $110+130 \times $135 + 75\times $145=$44025\]The total number of shares is\[50+110+130+75=365\]The mean cost per share to two decimal places is then,\[\frac{$44025}{365}=$120.62\]

OpenStudy (anonymous):

Thank you that was right

OpenStudy (anonymous):

welcome